# Topological properties of domain walls in antiferromagnetic topological insulators

### v1

@@ Abstract
Motivated by the study of stacking faults in weak topological insulators and the observation of magnetic domain walls in MnBi2nTe3n+1, we explore the topological properties of domain walls in antiferromagnetic topological insulators. We develop two tight-binding models: one based on a strong topological insulator with antiferromagnetic order, and another built from stacked Chern insulators with alternating Chern numbers. Both systems are dual topological insulators, i.e. they are at the same time antiferromagnetic and crystalline topological insulators, but differ by the type of mirror symmetry protecting the crystalline phase: spinful versus spinless. We show that in the spinful case the mirror Chern number is invariant under time reversal and that it changes sign in the spinless case. This influences the properties of the two systems in the presence of a magnetic domain wall, which is created in the system when the magnetization is flipped via a time-reversal transformation. In the first type, the bulk of the domain wall is gapped but the defect will host chiral edge states when it ends on an external ferromagnetic surface. In the second, due to the flip in the mirror Chern number, the domain wall is a two-dimensional embedded semimetal with 2D gapless states protected by mirror symmetry. Our results show that domain walls can be a source of non-trivial topology, allowing to generate and manipulate gapless states within the bulk and the ferromagnetic surfaces of antiferromagnetic topological insulators.
@@ I Introduction
We consider two different types of AFTI, which we label as type-I and type-II. The first (type-I) is made by adding antiferromagnetic order to an STI. A well-known AFTI in this class is MnBi2nTe3n+1 [5, 6]. This material is composed of a stack of MnBi2Te4 magnetic septuple layers with $n-1$ quintuple layers of Bi2Te3 between them. The magnetization of the Mn atoms orders antiferromagnetically in the stacking direction. Bulk MnBi2nTe3n+1 is an AFTI with gapless states on type A surfaces and with gapped F surfaces [5, 6]. However, the material can exhibit different topological phases depending on the geometry, making it an important platform for the study of magnetic topological phases. For example, the topology changes in the quasi 2D limit, when a stack consisting of only a few layers is considered. When the system has an odd number of layers, it is a Chern insulator [39, 40, 41, 42]. In the even layer case, the quasi-2D stack does not host a quantum Hall effect, but the so-called quantum layer-Hall effect arises in the presence of an external magnetic field [43, 44].
@@
Lattice defects, such as domain walls that break the $S$ symmetry [45] and step edges [46], have already been studied in this compound family. Interestingly, the latter hosts topological modes [46], due to the presence of a half-integer quantum Hall effect on the ferromagnetic surfaces of the material. Domain walls that do not break the $S$ symmetry have been observed in MnBi2Te4 [47] and their topological properties have been studied in stacks with an odd number of layers. In this case, the defect has been predicted to host chiral modes [48].
@@
The second type of AFTI that we study (type-II) is made up of a stack of weakly-coupled Chern insulating layers, with adjacent layers carrying opposite Chern numbers. This type of topological phase could be realized in van der Waals materials, e.g., in stacks of graphene layers with opposite out-of-plane magnetization [49]. When a mirror symmetry is present, both types of AFTI are so-called dual topological insulators (DTIs) [50], which means that they are at the same time AFTIs carrying a non-trivial $\mathbb{Z}_{2}$ invariant and TCIs carrying a non-trivial mirror Chern number [27, 28, 29, 49].
@@
In this work, we consider domain walls created by flipping the magnetization in part of the system via a time-reversal transformation. The topological properties of the defect will depend on the effect of time reversal on the mirror Chern number of the system. As we show, in type-I AFTIs time reversal does not change the mirror Chern number. In this case the bulk of the domain wall is gapped, but chiral edge states are present when the defect ends on an external F surface. In the type-II AFTI on the other hand, the mirror Chern number changes sign under time reversal. In this case, the domain wall hosts 2D gapless modes protected by mirror symmetry, since it separates two regions with opposite mirror Chern number. Our main findings are summarized in Fig. 1.
@@
Figure 1: We consider magnetic domain walls (vertical green plane) formed between two regions with opposite magnetization (blue/yellow) in AFTIs. In the type-I AFTI, the domain wall hosts chiral edge modes (red arrows), provided it ends on one of the ferromagnetic surfaces. In the type-II AFTI, the domain wall is an embedded topological semimetal hosting Dirac cones protected by mirror symmetry.
@@ II Antiferromagnetic topological insulator from strong topological phase
$\begin{split}\tilde{M}_{z}(k_{z}=0)=UM_{z}(0)U^{\dagger}=\begin{pmatrix}i\mathbb{I}_{4\times 4}&0\\ 0&-i\mathbb{I}_{4\times 4}\end{pmatrix}.\end{split}$ (10)
@@
This is due to the fact that the AFTI inherits the dual topology from the STI phase. In the time-reversal invariant case, for $M=0$, the system is an STI, so the (100) mirror-symmetric surface will host a surface Dirac cone at $k_{z}=0$. The states |$\ket{\Psi}$ and $\mathcal{T}\ket{\Psi}$ that cross at the Dirac point are time-reversal conjugate of each other, but at the same time they are also eigenstates of mirror symmetry. Since the operators $\mathcal{T}$ and $M_{z}(k_{z})$ commute, states that belong to the same Kramers pair have to carry opposite mirror eigenvalue $m=\pm i$, i.e.
@@ IV Invariance of mirror Chern number under time reversal
The presence of a non-trivial mirror Chern number has important consequences for the topological properties of the system in the presence of a domain wall, which we model by flipping the magnetization in part of the system by time reversal. If the time-reversal transformation does not change the sign of the mirror Chern number, the 2D domain wall remains gapped except at regions where it intersects the F surfaces. In that case chiral edge modes occur due to the half-integer quantum Hall effect of the type F surfaces. On the other hand, if the mirror Chern number is flipped by time reversal, the domain wall is an embedded semimetal [57] with 2D gapless states protected by mirror symmetry. To understand when the mirror Chern number is invariant under time reversal we can consider a generic TCI with Hamiltonian $H(k_{x},k_{y},k_{z})$ and with mirror symmetry:
@@
In this discussion, we consider a spinful time-reversal operator with $\mathcal{T}^{2}=-\mathbb{I}$ and we fix the gauge such that the mirror operator commutes with time reversal: $[\mathcal{T},M_{z}(k_{z}=0)]=0$. With this gauge choice, a spinful mirror symmetries have imaginary eigenvalues $m_{\pm}=\pm i$, whereas spinless mirror operators have real eigenvalues $m_{\pm}=\pm 1$.
@@
$\begin{split}&H^{\rm TR}(k_{x},k_{y},k_{z}=0)=\mathcal{T}H(-k_{x},-k_{y},0)\mathcal{T}^{-1}=\\ &\begin{pmatrix}\tau H_{m_{-}}(-k_{x},-k_{y})\tau^{-1}&0\\ 0&\tau H_{m_{+}}(-k_{x},-k_{y})\tau^{-1},\end{pmatrix},\end{split}$ (30)
@@
$-(U^{\phantom{{\dagger}}}_{\mathcal{T}}U_{\mathcal{T}}^{*})(\eta^{\phantom{{\dagger}}}_{\alpha}\eta^{\phantom{{\dagger}}}_{z}\eta_{\alpha}^{*})=\mathbb{I}_{N\times N}\eta^{\phantom{{\dagger}}}_{z}.$ (32)
In this case if we choose $U^{\phantom{{\dagger}}}_{\mathcal{T}}U_{\mathcal{T}}^{*}=\mathbb{I}_{N\times N}$, meaning that $\eta^{\phantom{{\dagger}}}_{\alpha}\eta_{\alpha}^{*}=-\eta_{0}\Rightarrow\eta_{\alpha}=\eta_{y}$, Eq. (32) would result in $-\eta_{y}\eta_{z}\eta^{*}_{y}=\eta_{z}$, which is false. The only possibility is then to choose $U^{\phantom{{\dagger}}}_{\mathcal{T}}U_{\mathcal{T}}^{*}=-\mathbb{I}_{N\times N}$, meaning $\eta_{\alpha}\eta_{\alpha}^{*}=\eta_{0}$ and rewrite the commutation condition as:
@@
$\begin{split}&H^{\rm TR}(k_{x},k_{y},k_{z}=0)=\mathcal{T}H(-k_{x},-k_{y},0)\mathcal{T}^{-1}=\\ &\begin{pmatrix}\tau H_{m_{+}}(-k_{x},-k_{y})\tau^{-1}&0\\ 0&\tau H_{m_{-}}(-k_{x},-k_{y})\tau^{-1},\end{pmatrix},\end{split}$ (34)
@@ V Domain walls: construction and topological properties
Figure 2: Band structure and LDOS of the type-I AFTI Hamiltonian in Eq. (2) in a ribbon geometry with two domain walls. The plots are obtained for $L_{x}=40$ and $L_{z}=20$ unit cells, $t=1$, $v=0.5$, $m=2$ and $M=1$. (a) Band structure of the system. The bands crossing the bulk gap (in blue) are associated with gapless states at the intersection between the domain wall and the top/bottom surfaces. (b) LDOS of the system at zero energy, showing the localization of the gapless modes, between $x=\frac{L_{x}}{2}-1$ and $x=\frac{L_{x}}{2}$ and between $x=0$ and $x=L_{x}-1$.
We can see the system as two topological insulators connected at the two domain walls, at left- and right-hand side of the defect. In this case the mirror symmetry protecting the TCI phase is spinful, which means that the time-reversal transformation does not flip the mirror Chern number of the topological insulator on the right-hand side. The crystalline topology therefore does not protect any topological state at the domain wall. Nonetheless, chiral edge states will appear due to the presence of a half-integer Hall effect on the F surfaces. If the domain wall were to be moved away from the top and bottom surfaces, for instance if the magnetization is only flipped for sites whose $z$ coordinates are away from the boundaries, then no gapless modes would be formed.
Using Kwant [58], a Python library for tight binding calculations, we solve numerically the tight binding model in Eq. (2) in the ribbon geometry. We find four topological degenerate bands that cross the bulk gap in the band structure of the system, see Fig. 2(a). By plotting the local density of state (LDOS):
$n(E,x,z)=\sum_{n,k_{y}}\delta\left(E-\epsilon_{n}(k_{y})\right)|\innerproduct{x,z}{u_{n}(k_{y})}|^{2}$ (35)
@@
In the type-II AFTI, the presence of dual topology modifies the topological properties of the defect. The domain wall hosts 2D gapless topological states protected by the spinless mirror symmetry. To show this, we construct the domain wall in a slab geometry with a finite size $L_{x}$ and two conserved momenta $k_{y}$ and $k_{z}$, since we are interested in studying the bulk of the defect. We flip the magnetization at every site $x\geq\frac{L_{x}}{2}$ in the model in Eq. (16) by applying time-reversal symmetry, which exchanges the layer Hamiltonians. We add periodic boundary conditions (PBC) in the $x$ direction, thus creating two domain walls, between $x=\frac{L_{x}}{2}-1$ and $x=\frac{L_{x}}{2}$ and between $x=0$ and $x=L_{x}-1$, as before.
Figure 3: Band structure and LDOS of the type-II AFTI Hamiltonian in Eq. (16) in a slab geometry with two domain walls. The plots where obtained for $L_{x}=40$, $t=v=1$, $u=1.5$ and $v_{z}=0.1$. (a) Band structure of the system on the $k_{z}=0$ plane. The blue bands are associated with gapless states localized at the domain walls. (b) LDOS of the system at zero energy which shows that the gapless states are localized at the two domain walls, between $x=\frac{L_{x}}{2}-1$ and $x=\frac{L_{x}}{2}$ and between $x=0$ and $x=L_{x}-1$.
When the domain wall is introduced in the type-II AFTI model, the two halves of the system, on the left and right sides of the defect at $x=\frac{L_{x}}{2}$, carry opposite mirror Chern numbers $+1$ and $-1$. This is due to the fact that the mirror Chern number is flipped by the time-reversal transformation. Thus, the domain wall hosts two Dirac cones protected by mirror symmetry. For the same reason, the second domain wall, formed across the periodic boundary of the system, also hosts two Dirac cones.
@@ VI Conclusions
In the first type of AFTIs, the bulk of the domain wall is gapped, since time reversal does not flip the mirror Chern number. Nonetheless, when the domain wall ends on a F surface, edge states are present at its boundaries, since the F surfaces of the AFTI carry a half-integer Chern number. In a type-II AFTI, the topological properties of the domain wall are modified by the presence of the non-zero mirror Chern number. This is due to the fact that time reversal changes sign to the mirror Chern number in TCIs with spinless mirror symmetry. As a result, a mirror-symmetric domain wall forms an embedded 2D semimetal with gapless topological states protected by mirror symmetry. Domain walls which break mirror symmetry, e.g. those with a different orientation or with a magnetization component perpendicular to the domain wall plane, will generically be gapped out. However, gapless topological states can still be present if the mirror symmetry is locally preserved. For example a spherical domain wall in the bulk of the type-II AFTI can host gapless topological states at its equator. Controlling the geometry of domain walls could thus enable the engineering of the location of topological gapless states within the bulk of these topological magnetic materials.

### v2

@@ Abstract
Motivated by the study of stacking faults in weak topological insulators and the observation of magnetic domain walls in MnBi2nTe3n+1, we explore the topological properties of magnetic domain walls in antiferromagnetic topological insulators. We develop two tight-binding models for two different types of antiferromagnetic topological insulators: the first type obtained by adding antiferromagnetic order to a strong topological insulator, and another built from stacked Chern insulating layers with alternating Chern numbers. Both systems are dual topological insulators, i.e. they are at the same time antiferromagnetic and crystalline topological insulators, but differ by the type of mirror symmetry protecting the crystalline phase: spinful versus spinless. We show that in the spinful case the mirror Chern number is invariant under time reversal and that it changes sign in the spinless case. This influences the properties of the two systems in the presence of a magnetic domain wall, which we model as an interface between two regions of opposite magnetization. In the first type, the bulk of the magnetic domain wall is gapped but the defect will host chiral edge states when it ends on an external ferromagnetic surface. In the second, due to the change in the sign of the mirror Chern number, the magnetic domain wall is a two-dimensional embedded semimetal with 2D gapless states protected by mirror symmetry. Our results show that magnetic domain walls can be a source of non-trivial topology, allowing to generate and manipulate gapless states within the bulk and the ferromagnetic surfaces of antiferromagnetic topological insulators.
@@ I Introduction
We consider two different types of AFTI, which we label as type-I and type-II. The first (type-I) is made by adding antiferromagnetic order to an STI. A well-known AFTI in this class is MnBi2nTe3n+1 [5, 6]. This material is composed of a stack of MnBi2Te4 magnetic septuple layers with n1n-1 quintuple layers of Bi2Te3 between them. The magnetization of the Mn atoms orders antiferromagnetically in the stacking direction. Bulk MnBi2nTe3n+1 is an AFTI with gapless states on type A surfaces and with gapped F surfaces [5, 6]. However, the material can exhibit different topological phases depending on the geometry, making it an important platform for the study of magnetic topological phases. For example, the topology changes in the quasi 2D limit, when a stack consisting of only a few layers is considered. When the system has an odd number of layers, it is a Chern insulator [39, 40, 41, 42]. In the even layer case, the quasi-2D stack is not in a Chern insulating phase, but the so-called quantum layer-Hall effect arises in the presence of an external magnetic field [43, 44].
@@
Defects, such as magnetic domain walls that break the $S$ symmetry [45] and step edges [46], have already been studied in this compound family. Interestingly, the latter host topological modes [46], due to the presence of a half-integer Chern number on the ferromagnetic surfaces of the material. Magnetic domain walls that do not break the $S$ symmetry have been observed in MnBi2Te4 [47, 48] and their topological properties have been studied in stacks with an odd number of layers [49] and at the intersection with the external ferromagnetic surfaces of the system [50, 51]. In these cases, the defect has been predicted to host chiral modes.
@@
The second type of AFTI that we study (type-II) is made up of a stack of weakly-coupled Chern insulating layers, with adjacent layers carrying opposite Chern numbers. This second type of topological phase does not appear in MnBi2nTe3n+1, but it could be realized in van der Waals materials, e.g., in stacks of graphene layers with opposite out-of-plane magnetization [52]. When a mirror symmetry is present, both types of AFTI are so-called dual topological insulators (DTIs) [53], which means that they are at the same time AFTIs carrying a non-trivial $\mathbb{Z}_{2}$ invariant and TCIs carrying a non-trivial mirror Chern number [27, 28, 29, 52].
In this work, we consider magnetic domain walls created by flipping the magnetization in part of the system. Thus, the domain wall is formed as an interface between two systems related to each other by a time-reversal transformation. Since magnetic domain walls are the only type of domain walls that we study in this work, the terms domain wall and magnetic domain wall are used interchangeably in the rest of the manuscript. Our main findings are summarized in Fig. 1. We are interested in studying the topological properties of the bulk of the defect. The domain wall is a 2D structure formed inside the 3D bulk of the material, shown in green in Fig. 1. We refer to the intersection between the 2D defect and the external surfaces of the system as boundaries of the domain wall, whereas we refer to the part of the defect buried in the bulk of the AFTI as bulk of the domain wall.
@@
The topological properties of the defect will depend on the effect of time reversal on the mirror Chern number of the system. As we show, in type-I AFTIs time reversal does not change the mirror Chern number. In this case the bulk of the domain wall is gapped, but chiral edge states, shown in red in Fig. 1, are present when the defect ends on an external F surface, similarly to Ref. [50, 51]. In the type-II AFTI on the other hand, the mirror Chern number changes sign under time reversal. In this case, the domain wall hosts 2D gapless modes protected by mirror symmetry, shown as a Dirac cone in Fig. 1, since it separates two regions with opposite mirror Chern number.
@@
Figure 1: We consider magnetic domain walls (vertical green plane) formed between two regions with opposite magnetization (blue/yellow) in AFTIs. The domain wall is a 2D structure that lies in the bulk of the 3D AFTIs and separates regions of the systems with opposite magnetization. We refer to the intersection between the domain wall and the external surfaces of the system as boundaries of the domain wall, whereas we refer to the part of the domain wall that lies away from the external surfaces as bulk of the 2D structure. In the type-I AFTI, the domain wall hosts chiral edge modes (red arrows), provided it ends on one of the ferromagnetic surfaces. In the type-II AFTI, the domain wall is an embedded topological semimetal hosting Dirac cones protected by mirror symmetry, with gapless states localized in the bulk of the defect (green plane).
@@ II Antiferromagnetic topological insulator from strong topological phase
[…]
@@
[…]
@@ IV Invariance of mirror Chern number under time reversal
The presence of a non-trivial mirror Chern number has important consequences for the topological properties of the system in the presence of a domain wall, which we model by flipping the magnetization in part of the system. As mentioned before, we construct the domain wall as an interface between two systems that are related to each other by a time reversal transformation. If the time-reversal transformation does not change the sign of the mirror Chern number, the 2D domain wall remains gapped except at regions where it intersects the F surfaces. In that case chiral edge modes occur due to the half-integer Chern number of the type F surfaces. On the other hand, if the mirror Chern number is flipped by time reversal, the domain wall is an embedded semimetal [60] with 2D gapless states protected by mirror symmetry. To understand when the mirror Chern number is invariant under time reversal we can consider a generic TCI with Hamiltonian $H(k_{x},k_{y},k_{z})$ and with mirror symmetry:
@@
In this discussion, we consider a spinful time-reversal operator with $\mathcal{T}^{2}=-\mathbb{I}$ and we fix the gauge such that the mirror operator commutes with time reversal: $\mathcal{T},M_{z}(k_{z}=0)]=0$. With this gauge choice, we say that the mirror symmetry is spinful if the mirror operator has imaginary eigenvalues $m_{\pm}=\pm i$, whereas the mirror symmetry is spinless if the eigenvalues are real $m_{\pm}=\pm 1$. Physically a spinful lattice symmetry is used to characterize systems with strong spin-orbit coupling, such as the type-I AFTIs, whereas systems with weak spin-orbit coupling, such as graphene, are characterized using a spinless symmetry.
@@
[…]
@@
$(U^{\phantom{{\dagger}}}_{\mathcal{T}}U_{\mathcal{T}}^{*})(\eta^{\phantom{{\dagger}}}_{\alpha}\eta^{\phantom{{\dagger}}}_{z}\eta_{\alpha}^{*})=\mathbb{I}_{N\times N}\eta^{\phantom{{\dagger}}}_{z}.$ (32)
In this case if we choose $U^{\phantom{{\dagger}}}_{\mathcal{T}}U_{\mathcal{T}}^{*}=\mathbb{I}_{N\times N}$, meaning that $\eta^{\phantom{{\dagger}}}_{\alpha}\eta_{\alpha}^{*}=-\eta_{0}\Rightarrow\eta_{\alpha}=\eta_{y}$, Eq. (32) would result in $\eta_{y}\eta_{z}\eta^{*}_{y}=\eta_{z}$, which is false. The only possibility is then to choose $U^{\phantom{{\dagger}}}_{\mathcal{T}}U_{\mathcal{T}}^{*}=-\mathbb{I}_{N\times N}$, meaning $\eta_{\alpha}\eta_{\alpha}^{*}=\eta_{0}$ and rewrite the commutation condition as:
@@
[…]
@@ V Domain walls: construction and topological properties
Figure 2: Band structure and LDOS of the type-I AFTI Hamiltonian in Eq. (2) in a ribbon geometry, with conserved momentum $k_{y}$ and finite size $L_{x}$ and $L_{z}$ in the $x$ and $z$ directions, with two domain walls. The plots are obtained for $L_{x}=40$ and $L_{z}=20$ unit cells, $t=1$, $v=0.5$, $m=2$ and $M=1$. All lengths are in units of the lattice constant $a=1$ and all energies are in units of the nearest neighbor hopping $t=1$. (a) Band structure of the system. The bands crossing the bulk gap (in blue) are associated with gapless states at the intersection between the domain wall and the top/bottom surfaces. (b) LDOS of the system at zero energy, showing the localization of the gapless modes, between $x=\frac{L_{x}}{2}-1$ and $x=\frac{L_{x}}{2}$ and between $x=0$ and $x=L_{x}-1$.
We can see the system as two topological insulators connected at the two domain walls, at left- and right-hand side of the defect. In this case the mirror symmetry protecting the TCI phase is spinful, which means that the time-reversal transformation does not flip the mirror Chern number of the topological insulator on the right-hand side. The crystalline topology therefore does not protect any topological state at the domain wall. Nonetheless, chiral edge states will appear due to the presence of a half-integer Chern number on the F surfaces [50, 51]. If the domain wall were to be moved away from the top and bottom surfaces, for instance if the magnetization is only flipped for sites whose $z$ coordinates are away from the boundaries, then no gapless modes would be formed.
Using Kwant [61], a Python library for tight binding calculations, we solve numerically the tight binding model in Eq. (2) in the ribbon geometry. Our code is available on Zenodo [62]. We find four topological degenerate bands that cross the bulk gap in the band structure of the system, see Fig. 2(a). By plotting the local density of state (LDOS):
[…]
@@
In the type-II AFTI, the presence of dual topology modifies the topological properties of the defect. The domain wall hosts 2D gapless topological states protected by the spinless mirror symmetry. To show this, we construct the domain wall in a slab geometry with a finite size $L_{x}$ and two conserved momenta $k_{y}$ and $k_{z}$, since we are interested in studying the bulk of the defect. Our code is available on Zenodo [62]. We flip the magnetization at every site $x\geq\frac{L_{x}}{2}$ in the model in Eq. (16) by applying time-reversal symmetry, which exchanges the layer Hamiltonians. We add periodic boundary conditions (PBC) in the $x$ direction, thus creating two domain walls, between $x=\frac{L_{x}}{2}-1$ and $x=\frac{L_{x}}{2}$ and between $x=0$ and $x=L_{x}-1$, as before.
Figure 3: Band structure and LDOS of the type-II AFTI Hamiltonian in Eq. (16) in a slab geometry, with a finite size $L_{x}$ and two conserved momenta $k_{y}$ and $k_{z}$, with two domain walls. The plots where obtained for $L_{x}=40$, $t=v=1$, $u=1.5$ and $v_{z}=0.1$. All lengths are in units of the lattice constant $a=1$ and all energies are in units of the nearest neighbor hopping $t=1$. (a) Band structure of the system on the $k_{z}=0$ plane. The blue bands are associated with gapless states localized at the domain walls. (b) LDOS of the system at zero energy which shows that the gapless states are localized at the two domain walls, between $x=\frac{L_{x}}{2}-1$ and $x=\frac{L_{x}}{2}$ and between $x=0$ and $x=L_{x}-1$.
When the domain wall is introduced in the type-II AFTI model, the two halves of the system, on the left and right sides of the defect at $x=\frac{L_{x}}{2}$, carry opposite mirror Chern numbers +1+1 and 1-1. This is due to the fact that the mirror Chern number is flipped by the time-reversal transformation. Thus, the domain wall hosts two Dirac cones protected by mirror symmetry. For the same reason, the second domain wall, formed across the periodic boundary of the system, also hosts two Dirac cones.
@@ VI Conclusions
In the first type of AFTIs, the bulk of the domain wall is gapped, since time reversal does not flip the mirror Chern number. Nonetheless, when the domain wall ends on a F surface, edge states are present at its boundaries, since the F surfaces of the AFTI carry a half-integer Chern number [50, 51]. In a type-II AFTI, the topological properties of the domain wall are modified by the presence of the non-zero mirror Chern number. This is due to the fact that time reversal changes sign to the mirror Chern number in TCIs with spinless mirror symmetry. As a result, a mirror-symmetric domain wall forms an embedded 2D semimetal with gapless topological states protected by mirror symmetry. Domain walls which break mirror symmetry, e.g. those with a different orientation or with a magnetization component perpendicular to the domain wall plane, will generically be gapped out. However, gapless topological states can still be present if the mirror symmetry is locally preserved. For example a spherical domain wall in the bulk of the type-II AFTI can host gapless topological states at its equator. Controlling the geometry of domain walls could thus enable the engineering of the location of topological gapless states within the bulk of these topological magnetic materials.